Calculate the standard cell potential at 25 ∘C for the reaction X(s)+2Y+(aq)→X2+(aq)+2Y(s) where ΔH∘ = -923 kJ and ΔS∘ = -257 J/K . Express your answer to three significant figures and include the appropriate units.
Answer #1

ΔHo = -923.0 KJ/mol

ΔSo = -257 J/mol.K

= -0.257 KJ/mol.K

T= 25.0 oC

= (25.0+273) K

= 298 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -923.0 - 298.0 * -0.257

ΔGo = -846.414 KJ/mol

Here number of electrons being transferred = 2

ΔGo = -846.414 KJ/mol

ΔGo = -846414 J/mol

F = 96500 C

use:

ΔGo = -n*F*E

-846414= -2*96500*E

E = 4.3856 V

Answer: 4.39 V